# The Population Resemblance Statistic: A Chi-Square Measure of Fit for Banking

### v3

@@ 3.1 Small sample behavior
In this subsection, we investigate the finite sample properties of the PSI and PRS, examining both scenarios where the current population $\bm{p}$ matches and departs from the reference population $\bm{p}_{0}$ used in model construction. Critically, $\bm{p}_{0}$ is assumed to be fixed and known throughout the paper. Define the scaled statistics $T_{n}=n\times\mathrm{PSI}$ and $Q_{n}=n\times\mathrm{PRS}$ where $\mathrm{PSI}$ and $\mathrm{PRS}$ are given in (1) and (3), respectively. Both of these are known to follow a (central) $\chi^{2}$ distribution with $B-1$ degrees of freedom under the assumption of no population shift, $\bm{p}=\bm{p}_{0}$, as the sample size increases. However, in smaller samples, their behavior can deviate significantly from asymptotic expectations, making it important to understand these differences for practical applications.
Despite the well-established asymptotic distribution of $T_{n}$ (Kullback, , 1978, Chapter 6), practitioners often rely on the thresholds of Lewis, (1994), which prescribe $\mathrm{PSI}<0.1$ to indicate acceptable population similarity and $\mathrm{PSI}\geq 0.25$ as a trigger for model reconstruction. These thresholds do not account for critical factors such as sample size or the number of categories, which can substantially affect the behavior and interpretation of the PSI in finite-sample settings.
To illustrate the limitations of these fixed thresholds, we conducted a simulation study examining the probability of mandating model reconstruction under two scenarios:
@@
Table 1: Estimated probabilities of mandating
reconstruction under fixed PSI thresholds.
@@
The simulation results reveal two critical issues with fixed PSI thresholds: For small samples ($n=50$), reconstruction is mandated too frequently, even under no shift. For large samples ($n=500$), reconstruction is rarely triggered, even with moderate shifts.
@@
Figure 1: Mean stability ratios for
$T_{n}$ (solid) and $Q_{n}$ (dashed)
Figure 2: Variance stability ratios for
$T_{n}$ (solid) and $Q_{n}$ (dashed)
The PRS ($Q_{n}$) exhibits mean stability even at small sample sizes and variance stability beyond $n=50$. In contrast, the PSI ($T_{n}$) shows substantial instability, with empirical moments deviating from asymptotic values by up to 8% (mean) and 30% (variance) in small samples. While close empirical and asymptotic moments do not guarantee convergence, large discrepancies signal lack of asymptotic convergence by the specific sample size.
@@ 3.2 Non-central chi-square limiting distribution
Having established the superior small-sample properties of the PRS under no population shift in comparison to the PSI, we now examine its behavior under population shift, where the current population $\bm{p}$ may differ from $\bm{p}_{0}$ by up to a specified tolerance $\delta$. Recall that $Q_{n}=n\times\mathrm{PRS}$ denotes the sample-size normalized PRS statistic.
Now, for $j=1,\ldots,B$, define category-specific deviations $\delta_{j}=p_{j}-p_{0j}$. We assume these deviations are small, specifically of the order $n^{-1/2}$; that is, $p_{j}-p_{0j}=O(n^{-1/2})$. Formally, this implies the existence of finite constants $\xi_{j}$ such that $\lim_{n\rightarrow\infty}n^{1/2}\delta_{j}=\xi_{j}.$ In this framework, the true current probabilities are written as $p_{j}=p_{0j}+\delta_{j}$.
To evaluate the asymptotic distribution of $Q_{n}$ under these deviations, define
$Z_{j}=\frac{\sqrt{n}(\hat{p}_{j}-p_{j})}{\sqrt{p_{j}}},\ j=1,\ldots,B.$
Rewriting $Q_{n}$ in terms of the $Z_{j}$ yields
$Q_{n}=\sum_{j=1}^{B}\left\{\sqrt{\frac{p_{j}}{p_{0j}}}\,Z_{j}+\frac{\sqrt{n}(p_{j}-p_{0j})}{\sqrt{p_{0j}}}\,\right\}^{2}.$
This formulation makes explicit how deviations from the reference distribution $\bm{p}_{0}$ influence the large-sample behavior of the statistic. Under multinomial sampling, the $Z_{j}$ are asymptotically standard normal. Combined with the assumption that $p_{j}-p_{0j}=O(n^{-1/2})$, we have $\mathrm{Var}(\sqrt{p_{j}/p_{0j}}\,Z_{j})=1+O(n^{-1/2})$. Consequently, $Q_{n}$ can be shown to converge to a non-central chi-square distribution with $B-1$ degrees of freedom and a finite non-centrality parameter. The explicit form of this non-centrality parameter in our context is given in Proposition 1; further details of this convergence can be found in Cressie and Read, (1984).
@@
Under the assumption that $p_{j}=p_{0j}+\delta_{j}=O(n^{-1/2})$, $Q_{n}$ converges in distribution as $n\to\infty$ to a non-central chi-square distribution with $B-1$ degrees of freedom,
[…]
where the non-centrality parameter $\lambda$ is given by
@@
Since the true values of $p_{j}$ are typically unknown, the non-centrality parameter $\lambda$ cannot be calculated. Additionally, care is needed when interpreting $\lambda$, as the formula suggests a dependence on the sample size $n$. However, under the assumption that $p_{j}-p_{0j}$ decreases at a rate of $n^{-1/2}$, this dependence is offset, ensuring that $\lambda$ remains well-defined in the asymptotic framework.
To address the challenge of $\lambda$ being unknown, we adopt a conservative approach by framing the problem in terms of the maximal non-centrality parameter under $\delta$-resemblance. This aligns with the concept of least favorable distributions, where test statistics are evaluated under the worst-case scenario within the null hypothesis, see Reinhardt, (1961). This approach allows us to construct robust decision-making critical values using the supremum of $\lambda$, defined as
@@
A crucial property of the non-central chi-square distribution underpins this framework. Specifically, the distribution satisfies a stochastic ordering property: for $\lambda\leq\lambda_{\mathrm{sup}}$ with fixed degrees of freedom $B-1$,
@@
where $\preceq_{\mathrm{st}}$ denotes stochastic dominance. This means that for random variables $X\sim\chi^{2}_{B-1}(\lambda_{\mathrm{sup}})$ and $Y\sim\chi^{2}_{B-1}(\lambda)$, the cumulative distribution functions satisfy $F_{Y}(x)\geq F_{X}(x)\text{ for all }x\in\mathbb{R}.$ This property ensures that the maximal non-centrality parameter $\lambda_{\mathrm{sup}}$ serves as a conservative basis for decision-making, enabling robust inference despite uncertainty about the true values of $\bm{p}$.
In the next proposition, we establish how the maximal non-centrality parameter $\lambda_{\mathrm{sup}}$ depends on the baseline probabilities $\bm{p}_{0}$ and the partition size $B$, providing a precise characterization of $\lambda_{\mathrm{sup}}$ in terms of these parameters.
@@
To derive the maximal non-centrality parameter $\lambda_{\mathrm{sup}}$, we consider the supremum of $\lambda$ under the constraint of $\delta$-resemblance, $\bm{p}\in\mathcal{P}(\delta\mid\bm{p}_{0})$. The convexity of $\lambda$ as a function of $\bm{p}$ ensures that its maximum is attained at an extreme point of the feasible set $\mathcal{P}(\delta\mid\bm{p}_{0})$. Translating $\mathcal{P}(\delta\mid\bm{p}_{0})$ by $\bm{p}_{0}$, this set becomes the intersection of a $\delta$-scaled $\ell_{\infty}$-ball (hypercube) and a co-dimension 1 subspace orthogonal to the main diagonal. Each extreme point of this set has coordinates $p_{j}\in\{p_{0j}-\delta,\,p_{0j},\,p_{0j}+\delta\}$, with at most one $p_{j}$ remaining at $p_{0j}$, and with the number of $+\delta$ and $-\delta$ deviations being equal. Two distinct cases need to be considered.
@@
Thus, the explicit form of $\lambda_{\mathrm{sup}}$ is established for both even and odd $B$, completing the proof. ∎
@@ 3.3 Decision-making framework for population monitoring
To implement the PRS for model monitoring, we propose a three-region decision-making framework that aligns with the expectations among risk management practitioners, as reflected by the Lewis constants for the PSI. This framework can, in principle, be extended to accommodate more than three regions, allowing for finer granularity in monitoring. It is parameterized by the risk tolerance $\delta>0$, representing the acceptable level of deviation from the model construction probabilities, and a multiplier $M>1$ such that $M\delta$ defines the boundary for full discrepancy (i.e., unacceptable level of deviation).
In addition, two decision sensitivity parameters, $\alpha_{1}$ and $\alpha_{2}$, control how readily the procedure transitions between the three decision regions. Specifically, $\alpha_{1}$ controls the sensitivity to model reconstruction, reflecting the likelihood of classifying a population as fully discrepant when it is still $\delta$-resemblant. Similarly, $\alpha_{2}$ controls the sensitivity to continued model use, indicating the likelihood of maintaining the current model when the population is at the boundary of $M\delta$-resemblance. Both serve as operational decision parameters, enabling practitioners to calibrate the framework based on their desired balance between stickiness [^1] The term “stickiness” is used here to denote the model’s resistance to change, which might otherwise be described as “stability.” However, we avoid “stability” to prevent confusion with population stability as used throughout this paper, where it has a distinct technical meaning. (the tendency to favor model continuity) and responsiveness (the ability to quickly detect and react to population shifts). These parameters determine the decision regions for the PRS framework as follows:
@@
The decision regions for the PRS framework are defined as follows, with $0\leq\alpha_{1},\alpha_{2}\leq 1$ corresponding to the level of risk aversion for erroneous decisions,
$\mathcal{R}_{1}$ =$=\left\{\mathrm{PRS}\leq\tau_{1}\right\}$ (Continue using model, “acceptable”)
$\mathcal{R}_{2}$ =$=\left\{\tau_{1}<\mathrm{PRS}\leq\tau_{2}\right\}$ (Enhanced monitoring, “partially discrepant”)
$\mathcal{R}_{3}$ =$=\left\{\mathrm{PRS}>\tau_{2}\right\}$ (Reconstruct model, “fully discrepant”)
where the region boundaries, or “critical values”, are
$\tau_{1}:=\frac{F_{B-1}^{-1}(\alpha_{2},\,M^{2}\lambda_{\mathrm{sup}})}{n}\ \ \ \text{and}\ \ \ \tau_{2}:=\frac{F_{B-1}^{-1}(1-\alpha_{1},\,\lambda_{\mathrm{sup}})}{n}.$
Herein, $F_{B-1}^{-1}(\alpha,\,\lambda)$ denotes the inverse cdf, i.e., the $100\alpha$-th percentile, of the non-central $\chi^{2}$-distribution with degrees of freedom $B-1$ and non-centrality parameter $\lambda$.
These decision regions are schematically illustrated in Figure 3 that depicts the PRS density curves derived from the non-central $\chi^{2}$ distribution. Tail regions corresponding to $\alpha_{1}$ and $\alpha_{2}$ are highlighted.
@@
Practitioners should carefully select parameters $\alpha_{1}$ and $\alpha_{2}$, as excessively large values may cause the intermediate monitoring region $\mathcal{R}_{2}$ to shrink or disappear entirely if the decision boundaries overlap. Our findings suggest that values in the range $[0.01,0.2]$ generally yield stable results, though a trade-off may be necessary if greater sensitivity is required in one direction or the other. Furthermore, setting excessively large $M$ can distort the framework by shifting the critical values too far apart, undermining the reliability of monitoring decisions.
As noted by a reviewer, the decision framework can be interpreted in a hypothesis-testing framework, where $H_{01}:\bm{p}\in\mathcal{P}(\delta\,|\,\bm{p}_{0})$ and $H_{02}:\bm{p}\in\mathcal{P}(M\delta\,|\,\bm{p}_{0})$ assert that the population is $\delta$-resemblant or $M\delta$-resemblant, respectively. Since $M>1$, the set of $\delta$-resemblant distributions is strictly contained within the set of $M\delta$-resemblant distributions, making $H_{01}$ nested within $H_{02}$. This structure enables a tiered decision-making process: rejecting $H_{01}$ but not $H_{02}$ places the model in the partially discrepant monitoring region $\mathcal{R}_{2}$, signaling a deviation beyond $\delta$ that does not yet exceed $M\delta$. This allows for increased scrutiny without immediate model reconstruction. In contrast, rejecting both hypotheses indicates a more substantial shift, necessitating full model reconstruction. The approach of using two null hypotheses for the same parameter(s) is uncommon but not new; for example, Schuirmann, (1987) applies a related framework in bioequivalence testing.
Within this framework, $\alpha_{1}$ serves as a Type I error rate under $H_{01}$, determining the likelihood of rejecting both $H_{01}$ and $H_{02}$ when the population is (only) $\delta$-resemblant. This corresponds to incorrectly concluding that the population is fully discrepant, leading to model reconstruction. As such, $\alpha_{1}$ defines the decision boundary for crossing from $\mathcal{R}_{2}$ to $\mathcal{R}_{3}$ and governs the corresponding upper critical value $\tau_{2}$ in Definition 2.
Conversely, $1-\alpha_{2}$ is the likelihood of exiting $\mathcal{R}_{1}$ when the population is at the boundary of $M\delta$-resemblance. While this resembles a detection of power, it is still defined within the null hypothesis framework of $H_{02}$ rather than, as conventional power, against an alternative hypothesis. Instead, $\alpha_{2}$ controls the likelihood of maintaining the current model despite a potential shift of magnitude $M\delta$. Thus, it defines the decision boundary for crossing from $\mathcal{R}_{1}$ to $\mathcal{R}_{2}$, i.e. the lower critical value $\tau_{1}$ in Definition 2.
This structure, informed by the stochastic dominance of the non-central chi-square distribution, upholds a conservative decision-making process while allowing practitioners to fine-tune decision sensitivity and maintain clear boundaries for monitoring and intervention.

### v4

@@ 3.1 Small sample behavior
In this subsection, we investigate the finite sample properties of the PSI and PRS, examining both scenarios where the current population $\bm{p}$ matches and departs from the reference population $\bm{p}_{0}$ used in model construction. Critically, $\bm{p}_{0}$ is assumed to be fixed and known throughout the paper. Define the scaled statistics $T_{n}=n\times\mathrm{PSI}$ and $Q_{n}=n\times\mathrm{PRS}$ where $\mathrm{PSI}$ and $\mathrm{PRS}$ are given in (1) and (4), respectively. Both of these are known to follow a (central) $\chi^{2}$ distribution with B1B-1 degrees of freedom under the assumption of no population shift, $\bm{p}=\bm{p}_{0}$, as the sample size increases. However, in smaller samples, their behavior can deviate significantly from asymptotic expectations, making it important to understand these differences for practical applications.
Despite the well-established asymptotic distribution of $T_{n}$ [kullback1978information, Chapter 6], practitioners often rely on the thresholds of [lewis1994introduction], which prescribe $\mathrm{PSI}<0.1$ to indicate acceptable population similarity and $\mathrm{PSI}\geq 0.25$ as a trigger for model reconstruction. These thresholds do not account for critical factors such as sample size or the number of categories, which can substantially affect the behavior and interpretation of the PSI in finite-sample settings.
To illustrate the limitations of fixed thresholds, we conducted a simulation study examining the probability of mandating model reconstruction under two scenarios:
@@
The simulation results reveal two critical issues with fixed PSI thresholds: For small samples ($n=50$), reconstruction is mandated too frequently, even under no shift. For larger samples ($n=500$), reconstruction is rarely triggered, even with moderate shifts.
Table 1: Estimated probabilities of mandating reconstruction under fixed PSI thresholds.
@@
Figure 1: Mean stability ratios for $T_{n}$ (solid) and $Q_{n}$ (dashed).
Figure 2: Variance stability ratios for $T_{n}$ (solid) and $Q_{n}$ (dashed).
After appropriate adjustment for sample size, the PRS (as per $Q_{n}$) exhibits mean stability even at small sample sizes and variance stability for $n\geq 50$. In contrast, the PSI (as per $T_{n}$) shows substantial instability, with empirical moments deviating from asymptotic values by up to 8% (mean) and 30% (variance) in small samples. Although close agreement between empirical and asymptotic moments does not ensure convergence, substantial discrepancies indicate a lack of asymptotic convergence at the corresponding sample size.
@@ 3.2 Non-central chi-square limiting distribution
Having established the superior small-sample properties of the PRS under no population shift in comparison to the PSI, we now examine PRS behavior under population shift, where the current population $\bm{p}$ may differ from the reference population $\bm{p}_{0}$ by up to a specified tolerance $\delta$. Recall the sample-size normalized PRS statistic,
$Q_{n}=n\times\mathrm{PRS}=n\sum_{j=1}^{B}\frac{(\hat{p}_{j}-p_{0j})^{2}}{p_{0j}}.$
Let $\delta_{j}=p_{j}-p_{0j}$ denote the category-specific deviations, assumed to be small and of the order $n^{-1/2}$; that is, $p_{j}-p_{0j}=\text{O}(n^{-1/2})$. Formally, we assume there exist constants $\xi_{j}$ such that $\lim_{n\rightarrow\infty}n^{1/2}\delta_{j}=\xi_{j}.$ Thus, we write $p_{j}=p_{0j}+\delta_{j}$, where $\bm{p}$ is a so-called local alternative near $\bm{p}_{0}$.
Under multinomial sampling, the count vector $(n_{1},\ldots,n_{B})^{\top}=n\times(\hat{p}_{1},\ldots,\hat{p}_{B})^{\top}$ follows a $\text{Multinomial}(n,\bm{p})$ distribution. By the multivariate central limit theorem for proportions, we have
$\sqrt{n}(\widehat{\bm{p}}-\bm{p})\xrightarrow{d}\text{N}\big(0,\bm{\Sigma}(\bm{p})\big),$
where $\bm{\Sigma}(\bm{p})=\mathrm{diag}(\bm{p})-\bm{p}\bm{p}^{\top}$ is a singular $B\times B$ covariance matrix of rank B1B-1, reflecting the constraint that the probabilities sum to one.
Expanding $Q_{n}$ using the decomposition $\hat{p}_{j}-p_{0j}=(\hat{p}_{j}-p_{j})+(p_{j}-p_{0j})$, and letting $\bm{\delta}=\bm{p}-\bm{p}_{0}$ and $\mathbf{W}=\mathrm{diag}(\bm{p}_{0})$, gives
$Q_{n}$ $=$ $\sum_{j=1}^{B}\left[\frac{n(\hat{p}_{j}-p_{j})^{2}}{p_{0j}}+\frac{2n(\hat{p}_{j}-p_{j})(p_{j}-p_{0j})}{p_{0j}}+\frac{n(p_{j}-p_{0j})^{2}}{p_{0j}}\right]$
$=$ $n(\widehat{\bm{p}}-\bm{p})^{\top}\mathbf{W}^{-1}(\widehat{\bm{p}}-\bm{p})+2n\bm{\delta}^{\top}\mathbf{W}^{-1}(\widehat{\bm{p}}-\bm{p})+n\bm{\delta}^{\top}\mathbf{W}^{-1}\bm{\delta}.$
Above, the first term converges to a chi-square distribution with B1B-1 degrees of freedom, $n(\widehat{\bm{p}}-\bm{p})^{\top}\mathbf{W}^{-1}(\widehat{\bm{p}}-\bm{p})\xrightarrow{d}\chi^{2}_{B-1}.$ The second term is linear in the centered random vector and satisfies $2n\bm{\delta}^{\top}\mathbf{W}^{-1}(\widehat{\bm{p}}-\bm{p})=2\bm{\xi}^{\top}\mathbf{W}^{-1}\sqrt{n}(\widehat{\bm{p}}-\bm{p})+\text{o}_{p}(1),$ which converges in distribution to a normal random variable. Here, $\text{o}_{p}(1)$ denotes a remainder term that converges to zero in probability. The third term is deterministic and forms the non-centrality parameter, $\lambda=n\bm{\delta}^{\top}\mathbf{W}^{-1}\bm{\delta}=\sum_{j=1}^{B}n\delta_{j}^{2}/p_{0j}.$ Combining these results and applying standard theory on quadratic forms under local alternatives ([cressie1984multinomial]), we obtain the following limiting distribution.
@@
Assume $\delta_{j}=p_{j}-p_{0j}=\text{O}(n^{-1/2})$ for $j=1,\ldots,B$. Then, as $n\to\infty$,
$Q_{n}\overset{d}{\longrightarrow}\chi^{2}_{B-1}(\lambda),$
a non-central chi-square distribution with B1B-1 degrees of freedom and non-centrality parameter $\lambda$ given by
@@
Since the true values of $p_{j}$ are typically unknown, direct evaluation of $\lambda$ is infeasible. Moreover, while the definition of $\lambda$ suggests a dependence on sample size $n$, it remains well-defined in the asymptotic framework due to the assumed $n^{-1/2}$ decay in the $\delta_{j}$.
To address the challenge of $\lambda$ being unknown, we adopt a conservative approach by framing the problem in terms of the maximal non-centrality parameter under $\delta$-resemblance. This aligns with the concept of least favorable distributions, where test statistics are evaluated under the worst-case scenario within the null hypothesis, see [reinhardt1961use]. This approach allows us to construct robust decision-making critical values using the supremum of $\lambda$, defined as
@@
A key property of the non-central chi-square distribution underpins this framework: stochastic ordering. For $\lambda\leq\lambda_{\mathrm{sup}}$ with fixed degrees of freedom B1B-1,
@@
where $\preceq_{\mathrm{st}}$ denotes stochastic dominance. That is, if $X\sim\chi^{2}_{B-1}(\lambda_{\mathrm{sup}})$ and $Y\sim\chi^{2}_{B-1}(\lambda)$, the cumulative distribution functions satisfy $F_{Y}(x)\geq F_{X}(x)\text{ for all }x\in\mathbb{R}.$ This property ensures that the maximal non-centrality parameter $\lambda_{\mathrm{sup}}$ provides a conservative basis for decision-making, enabling robust inference despite uncertainty about the true $\bm{p}$.
In the next proposition, we establish how the maximal non-centrality parameter $\lambda_{\mathrm{sup}}$ depends on the reference probabilities $\bm{p}_{0}$ and the partition size $B$, providing a precise characterization of $\lambda_{\mathrm{sup}}$ in terms of these parameters.
@@
To derive the maximal non-centrality parameter $\lambda_{\mathrm{sup}}$, we consider the supremum of $\lambda$ under the constraint of $\delta$-resemblance, $\bm{p}\in\mathcal{P}(\delta\,|\,\bm{p}_{0})$. The convexity of $\lambda$ as a function of $\bm{p}$ ensures that its maximum is attained at an extreme point of the feasible set $\mathcal{P}(\delta\,|\,\bm{p}_{0})$. Translating $\mathcal{P}(\delta\,|\,\bm{p}_{0})$ by $\bm{p}_{0}$, this set becomes the intersection of a $\delta$-scaled Chebyshev ball (i.e., a hypercube under the $\ell_{\infty}$ norm) and a codimension-1 vector subspace of $\mathbb{R}^{B}$ that passes through the center of the hypercube and is orthogonal to the main diagonal. Each extreme point of this set has coordinates $p_{j}\in\{p_{0j}-\delta,\,p_{0j},\,p_{0j}+\delta\}$, with at most one $p_{j}$ remaining at $p_{0j}$, and with the number of $\delta$ and $\delta$ deviations being equal. Two distinct cases need to be considered.
@@
Thus, $\lambda_{\mathrm{sup}}$ is now fully characterized for all $B$, completing the proof. ∎
@@ 3.3 Decision-making framework for population monitoring
As previewed in Section 1.1, the PRS framework replaces the traditional point null with a pair of nested composite hypotheses, enabling a three-tiered response structure that distinguishes tolerable from material shifts and supports practical risk management. This section formalizes the hypotheses, defines decision regions, and explains the operational role of tuning parameters.
With thanks to a reviewer for identifying this connection, we formalize our initial proposal using a nested hypothesis-testing framework based on the concept of $\delta$-resemblance. Let $H_{01}:\bm{p}\in\mathcal{P}(\delta\mid\bm{p}_{0})$ denote the hypothesis that the current population resembles the reference population $\bm{p}_{0}$ up to a maximum deviation of $\delta$, as defined in Section 2. Let $H_{02}:\bm{p}\in\mathcal{P}(M\delta\mid\bm{p}_{0})$ denote a more relaxed hypothesis allowing deviation up to $M\delta$, with $M>1$ a user-specified multiplier. These hypotheses are nested, with $H_{01}\subset H_{02}$.
This structure supports a three-level decision process, determined by which hypotheses are rejected. If neither $H_{01}$ nor $H_{02}$ is rejected, the model is considered acceptable for continued use. If $H_{01}$ is rejected but $H_{02}$ is not, the population shift exceeds the threshold of $\delta$ but remains within the bounds of $M\delta$, prompting enhanced monitoring. If both hypotheses are rejected, the shift is deemed substantial enough to warrant full model reconstruction.
@@
The PRS decision regions are defined as follows,
$\mathcal{R}_{1}$ $=\left\{\mathrm{PRS}\leq\tau_{1}\right\}$ (Continue using model, “acceptable”)
$\mathcal{R}_{2}$ $=\left\{\tau_{1}<\mathrm{PRS}\leq\tau_{2}\right\}$ (Enhanced monitoring, “partially discrepant”)
$\mathcal{R}_{3}$ $=\left\{\mathrm{PRS}>\tau_{2}\right\}$ (Reconstruct model, “fully discrepant”).
These decision boundaries, or “critical values”, are defined as
$\tau_{1}:=\frac{F_{B-1}^{-1}(\alpha_{2},\,M^{2}\lambda_{\mathrm{sup}})}{n}\ \ \ \text{and}\ \ \ \tau_{2}:=\frac{F_{B-1}^{-1}(1-\alpha_{1},\,\lambda_{\mathrm{sup}})}{n},$
where $0\leq\alpha_{1},\alpha_{2}\leq 1$ reflect the organization’s risk tolerance for erroneous decisions, and $F_{B-1}^{-1}(\alpha,\lambda)$ denotes the $\alpha$-quantile of the non-central $\chi^{2}$ distribution with B1B-1 degrees of freedom and non-centrality parameter $\lambda$.
The labels “acceptable”, “partially” and “fully discrepant” are illustrative and not industry agnostic. In banking, for example, a red-amber-green ($\mathbbm{RAG}$) classification is commonly used to describe degrees of discrepancy. Then, $\mathbbm{R}$ indicates membership to $\mathcal{R}_{3}$, $\mathbbm{A}$ to $\mathcal{R}_{2}$ and $\mathbbm{G}$ to $\mathcal{R}_{1}$. These descriptive labels should be tailored to the application context. Alternatives might include “low, medium, high risk of deviance.”
Within this framework, $\alpha_{1}$ serves as a Type I error rate under $H_{01}$, determining the likelihood of rejecting both $H_{01}$ and $H_{02}$ when the population is (only) $\delta$-resemblant. This corresponds to incorrectly concluding that the population is fully discrepant, leading to model reconstruction. As such, $\alpha_{1}$ defines the decision boundary for crossing from $\mathcal{R}_{2}$ to $\mathcal{R}_{3}$ and governs the corresponding upper critical value $\tau_{2}$ in Definition 2. Conversely, $1-\alpha_{2}$ is the likelihood of exiting $\mathcal{R}_{1}$ when the population is at the boundary of $M\delta$-resemblance. While this resembles a detection of power, it is still defined within the null hypothesis framework of $H_{02}$ rather than, as conventional power, against an alternative hypothesis. Instead, $\alpha_{2}$ controls the likelihood of maintaining the current model despite a potential shift of size $M\delta$. Thus, it defines the decision boundary for crossing from $\mathcal{R}_{1}$ to $\mathcal{R}_{2}$ — the lower critical value $\tau_{1}$ in Definition 2. The decision regions are schematically illustrated in Figure 3 that depicts the PRS density curves derived from the non-central $\chi^{2}$ distribution. Tail regions corresponding to $\alpha_{1}$ and $\alpha_{2}$ are highlighted.
@@
Using two null hypotheses for the same parameter is rare but not unprecedented; [schuirmann1987comparison], for example, applies a related strategy in bioequivalence testing. The use of two nested composite null hypotheses is what permits $\alpha_{1}$ and $\alpha_{2}$ to be specified independently, each tied to a distinct decision boundary with separate practical implications. Unlike traditional approaches that apply two significance levels to a single null hypothesis, the nested structure separates the risks of overreaction (rejecting a tolerable deviation) from underreaction (failing to act on a material one). This separation supports more meaningful discussions with stakeholders, as the two types of monitoring errors can be aligned with different business consequences and risk tolerances. Moreover, this structure, informed by the stochastic dominance of the non-central chi-square distribution, upholds a conservative decision-making process while allowing practitioners to fine-tune decision sensitivity and maintain clear boundaries for monitoring and intervention.
While the stochastic ordering result in Section 3.2 ensures robust Type I error control, it does not extend to settings where $\bm{p}\notin\mathcal{P}(M\delta\mid\bm{p}_{0})$. In such cases, evaluating $\mathbb{P}(\mathrm{PRS}\in\mathcal{R}_{j})$ for $j=1,2,3$ requires either specifying a concrete alternative or working within a least favorable alternative framework. These approaches introduce stronger modeling assumptions and greater technical complexity, which fall outside the scope of this paper’s focus on error calibration. Developing such power-based refinements is an important direction for future research.